OpenStudy (anonymous):

For any nonnegative real number c, = (root c)^2 A. c B. root c C. c2 D. 1

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle c=\left(\sqrt{c}\right)^2 }\) is this your statement?

OpenStudy (solomonzelman):

Or is your equation, \(\large\color{#000000 }{ \displaystyle \left(\sqrt{c}\right)^2 }\), and you are just trying to figure out what result this is going to yeild?